# Optimal and Efficient Sample Size Re-estimation: A Dynamic Cost Framework

### v1

@@ 2 Background of the JT Framework
We consider a clinical trial with a continuous endpoint, where measurements for the treatment and control arms are denoted by $Y_{i}^{trt}$ and $Y_{i}^{ctl}$, respectively. Assuming normality with a common known variance $\sigma^{2}$, we have:
@@
The treatment effect $\theta$ is defined as the difference in means $\theta=\mu_{t}-\mu_{c}$. An interim analysis is planned after collecting $n_{1}$ measurements. The observed treatment effect is $\hat{\theta}_{1}=\bar{Y}^{trt}-\bar{Y}^{ctl}$ and the corresponding standardized test statistic is denoted by $Z_{1}$. The conditional power, given $Z_{1}=z_{1}=\hat{\theta}_{1}/\sqrt{4\sigma^{2}/n_{1}}$ and assuming the true treatment effect is $\theta$, is defined as:
$CP_{\theta}(z_{1})=P_{\theta}\left(Z_{2}>1.96|Z_{1}=z_{1}\right)\,$, (2.1)
where $Z_{2}$ is the final standardized test statistic.
@@
$n_{2}^{JT}(z_{1})=\arg\max_{n_{2}(z_{1})}\left\{CP_{{\theta}}(z_{1},n_{2}(z_{1}))-\gamma(n_{2}(z_{1})-n_{min})\right\}\,$, (2.2)
where $n_{min}$ is the minimum total sample size (e.g., the original planned sample size), and the maximization is performed over the set of allowable sample sizes, $n_{2}(z_{1})$. The term $CP_{{\theta}}(z_{1},n_{2}(z_{1}))$ represents the conditional power, given $Z_{1}=z_{1}$, evaluated under a prespecified effect size of interest, $\theta$}. The quantity $\gamma$ serves as a tuning parameter that represents the acceptable “rate of exchange” between conditional power and sample size. It can be tied to the cost of additional subjects and is used to control the extent to which the sample size may be increased based on the interim data.
@@ 3 Optimal Sample Size Re-estimation with Dynamic Cost Functions
$\arg\max_{n_{2}(z_{1})}\left\{CP_{\theta}(z_{1},n_{2}(z_{1}))-\gamma(z_{1})n_{2}(z_{1})\right\}\,.$ (3.1)
@@ 3.1 Likelihood-ratio based $\gamma(z_{1})$
$\max_{n_{2}(z_{1})}P_{\theta}(\text{Rej }H_{0}|\Theta_{1})\hskip 14.22636pt\text{subject to}\hskip 14.22636ptE_{0}[n_{2}(z_{1})]\leq B_{0},\hskip 14.22636pt\text{and}\hskip 14.22636ptE_{\theta}[n_{2}(z_{1})]\leq B_{\theta}\,$, (3.2)
@@
The Lagrangian for this problem, with non-negative Lagrange multipliers $\lambda_{1}$ and $\lambda_{2}$, is given by:
$\mathcal{L}(n_{2}(z_{1}),\lambda_{1},\lambda_{2})=\int\left\{CP_{\theta}(z_{1},n_{2}(z_{1}))f_{\theta}(z_{1})-\left(\lambda_{1}f_{0}(z_{1})+\lambda_{2}f_{\theta}(z_{1})\right)n_{2}(z_{1})\right\}dz_{1}$
To maximize the Lagrangian, we maximize the integrand pointwise. Dividing the integrand by the density function $f_{\theta}(z_{1})$ transforms the problem into a pointwise maximization:
$\max_{n_{2}(z_{1})}\left\{CP_{\theta}(z_{1},n_{2}(z_{1}))-\left(\lambda_{2}+\lambda_{1}\frac{f_{0}(z_{1})}{f_{\theta}(z_{1})}\right)n_{2}(z_{1})\right\}\,.$ (3.3)
We denote the solution of this optimization problem by $n_{2}^{LR}(z_{1})$. This formulation yields a dynamic cost function in the JT fashion, where the cost is given by $\gamma(z_{1})=\lambda_{2}+\lambda_{1}\frac{f_{0}(z_{1})}{f_{\theta}(z_{1})}$. The term $\frac{f_{0}(z_{1})}{f_{\theta}(z_{1})}$ is the likelihood ratio. A large value of this ratio indicates that the observed interim statistic $z_{1}$ is more likely to arise under the null hypothesis, thereby increasing the cost of adding a sample and making it harder to justify a sample size increase.
3.1.1 Comparing $n_{2}^{JT}(z_{1})$ with $n_{2}^{LR}(z_{1})$
We now formally compare the performance of the JT rule, $n_{2}^{JT}(z_{1})$, with our proposed likelihood-ratio based rule, $n_{2}^{LR}(z_{1})$. For a fixed cost parameter $\gamma$, let $n_{2}^{JT}(z_{1},\gamma)$ be the optimal rule from the JT method. We claim that $n_{2}^{JT}(z_{1},\gamma)$ is not the optimal solution for the following constrained optimization problem:
$\min_{n_{2}(z_{1})}E_{0}[n_{2}(z_{1})]\quad\text{subject to}\quad P_{\theta}(\text{Rej }H_{0};n_{2}(z_{1}))=P_{\theta}(\text{Rej }H_{0};n_{2}^{JT}(z_{1}))\quad\text{and}\quad E_{\theta}[n_{2}(z_{1})]=E_{\theta}[n_{2}^{JT}(z_{1})]\,.$
@@ 3.1.1 Comparing $n_{2}^{JT}(z_{1})$ with $n_{2}^{LR}(z_{1})$
$\mathcal{L}(n_{2}(z_{1}),\mu_{1},\mu_{2})$ $=E_{0}[n_{2}(z_{1})]-\mu_{1}\left(P_{\theta}(\text{Rej }H_{0};n_{2}(z_{1}))-P_{\theta}(\text{Rej }H_{0};n_{2}^{JT}(z_{1}))\right)$
@@
For a function $n_{2}(z_{1})$ to be a minimizer of $\mathcal{L}(n_{2}(z_{1}),\mu_{1},\mu_{2})$, the integrand must be minimized for each $z_{1}$. We treat $n_{2}(z_{1})$ as a continuous variable for the purpose of optimization, a standard approach for this class of problems. While the sample size must be an integer, treating it as a continuous variable allows for the use of calculus-based optimization. The resulting continuous solution can then be rounded to the nearest integer. The performance loss from this approximation is typically negligible. Thus, a necessary condition for $n_{2}(z_{1})$ to be a minimizer is:
@@
For the JT method, the corresponding condition is $\frac{\partial CP_{\theta}(z_{1},n_{2}(z_{1}))}{\partial n_{2}(z_{1})}=\gamma$, which is a constant. Since the right-hand side of our derived condition is a function of $z_{1}$, this proves that $n_{2}^{JT}(z_{1})$ is not the rule that minimizes $E_{0}[n_{2}(z_{1})]$ for its given power and average sample size under the alternative hypothesis.
@@
3.1.2 Pareto optimality of $n_{2}^{LR}(z_{1})$
The likelihood-ratio based re-estimation rule, $n_{2}^{LR}(z_{1})$ is a solution to multi-objective optimization problem. As such, it possesses a fundamental property known as Pareto optimality.
@@ 3.1.2 Pareto optimality of $n_{2}^{LR}(z_{1})$
A rule $n_{2}^{*}(z_{1})$ is Pareto optimal if no other rule $n_{2}(z_{1})$ exists that is strictly better on at least one objective while being no worse on the others.
@@
For $\lambda_{1}\geq 0$ and $\lambda_{2}\geq 0$, $n_{2}^{LR}(z_{1})$ maximizes $\mathcal{L}(n_{2}(z_{1}))=P_{\theta}(\text{Rej }H_{0};n_{2}(z_{1}))-\lambda_{2}1E_{0}[n_{2}(z_{1})]-\lambda_{2}E_{\theta}[n_{2}(z_{1})]$. Assume that $n_{2}^{LR}(z_{1})$ is not Pareto optimal. By definition, this means there exists a rule $n^{*}_{2}(z_{1})$ such that
  1. $P_{\theta}(\text{Rej }H_{0};n_{2}^{*}(z_{1}))>P_{\theta}(\text{Rej }H_{0};n^{LR}_{2}(z_{1}))$  ,
@@ 2. $E_{0}[n^{*}_{2}(z_{1})]\leq E_{0}[n^{LR}_{2}(z_{1})]$  ,
Then, we got $\mathcal{L}(n^{*}_{2}(z_{1}))>\mathcal{L}(n^{LR}_{2}(z_{1}))$, which contradicts to the fact that $n_{2}^{LR}(z_{1})$ maximizes $\mathcal{L}(n_{2}(z_{1}))$.
The Pareto optimality of $n_{2}^{LR}(z_{1})$ establishes it as an efficient and well-balanced solution for the multi-objective problem of maximizing power while controlling the average sample size under both $\Theta_{0}$ and $\Theta_{1}$.
In the context of sample size re-estimation, Pareto optimality means that the likelihood-ratio based rule avoids wasteful designs. If one attempts to reduce sample size under the null, the only way to do so is by lowering power; conversely, if one increases power, additional participants must be recruited. Thus, $n_{2}^{LR}(z_{1})$ represents an efficiency frontier: it offers the best achievable balance between power and resource use. This provides reassurance to practitioners that adopting $n_{2}^{LR}(z_{1})$ does not leave potential efficiency gains unexploited.
3.1.3 Calibration of $\lambda_{1}$ and $\lambda_{2}$
To implement the likelihood-ratio based re-estimation rule, we must determine the values of the Lagrange multipliers $\lambda_{1}$ and $\lambda_{2}$. In this section, we propose a way that calibrating them to satisfy two specific constraints on the study design: the average sample size under the null hypothesis ($\Theta_{0}$) and the unconditional power under the alternative hypothesis ($\Theta_{1}$).
We aim to solve for a pair of $(\lambda_{1},\lambda_{2})$ that satisfies the following system of equations:
  1. 1.
    $g_{1}(\lambda_{1},\lambda_{2})=E_{0}[n_{2}^{LR}(\lambda_{1},\lambda_{2})]-B_{0}=0$
  2. 2.
    $g_{2}(\lambda_{1},\lambda_{2})=P_{\theta}(\text{Rej }H_{0};n_{2}^{LR}(\lambda_{1},\lambda_{2}))-\pi=0$
Here, $n_{2}^{LR}(\lambda_{1},\lambda_{2})$ is the optimal sample size re-estimation rule, $B_{0}$ is the target average sample size under $\Theta_{0}$, and $\pi$ is the target unconditional power under $\Theta_{1}$. We assume that $n_{2}\leq n_{2}^{LR}(z_{1})\leq N_{\max}$ for some maximum sample size $N_{\max}$. We also assume that $B_{0}<N_{\max}$ and $\pi<P_{\theta}(\text{Rej }H_{0};N_{\max})$, ensuring that the targets are achievable and do not require the maximum possible sample size.
We can establish the existence and uniqueness of the solution by analyzing the properties of the functions $g_{1}$ and $g_{2}$. The optimal sample size $n_{2}^{LR}(\lambda_{1},\lambda_{2})$ is a strictly decreasing function of both $\lambda_{1}$ and $\lambda_{2}$. Consequently, the average sample size under $\Theta_{0}$, $E_{0}[n_{2}^{LR}(\lambda_{1},\lambda_{2})]$, and the unconditional power, $P_{\theta}(\text{Rej }H_{0};n_{2}^{LR}(\lambda_{1},\lambda_{2}))$, are also strictly decreasing in both $\lambda_{1}$ and $\lambda_{2}$. This monotonic behavior is crucial.
Consider the boundary conditions for $\lambda_{1}$ and $\lambda_{2}$:
  • As $\lambda_{1},\lambda_{2}\to 0$: The cost of adding samples approaches zero, causing the optimal sample size to increase toward its maximum limit, $N_{\max}$. This implies that $g_{1}(\lambda_{1},\lambda_{2})\to N_{\max}-B_{0}>0$ and $g_{2}(\lambda_{1},\lambda_{2})\to P_{\theta}(\text{Rej }H_{0};N_{\max})-\pi>0$.
  • As $\lambda_{1},\lambda_{2}\to\infty$: The cost of adding samples becomes prohibitive. The optimal sample size decreases to its minimum value (i.e., the original planned sample size), causing both $g_{1}$ and $g_{2}$ to converge to a non-positive quantity.
Given the strict monotonicity and continuity of $g_{1}$ and $g_{2}$, the Intermediate Value Theorem guarantees the existence of a unique pair $(\lambda_{1},\lambda_{2})$ that solves the system of equations. Therefore, a root-finding algorithm, such as a multivariate Newton-Raphson method, can be used to numerically find the solution.
@@ 3.2 Return on investiment (ROI) based $\gamma(z_{1})$
Let $p(z_{1})=P(\text{effective}|z_{1})$ be the posterior probability of the treatment being effective, and $CP_{\theta}(z_{1},n_{2}^{ROI}(z_{1}))$ be the conditional power. The expected net return can then be expressed as:
@@ 4 Representation Theorem for Sample Size Re-estimation Rules
If the conditional power $CP_{\theta}(z_{1},n_{2})$ is differentiable and strictly concave with respective to $n_{2}$ for $z_{1}\in\mathcal{P}$, then for a SSR rule that yields an interior solution $n_{2}^{*}(z_{1}):\mathcal{P}\to(n_{min},N_{max})$, there exists a unique dynamic cost function $\gamma^{*}(z_{1})$ such that $n_{2}^{*}(z_{1})$ is the solution of the optimization problem:
$\max_{n_{2}\in(n_{min},N_{max})}\left\{CP_{\theta}(z_{1},n_{2})-\gamma^{*}(z_{1})n_{2}\right\}\,.$
@@
Let $n_{2}^{*}(z_{1})$ be an arbitrary, pre-specified SSR rule. Then, we can define a dynamic cost function, $\gamma^{*}(z_{1})$, as follows:
$\gamma^{*}(z_{1}):=\frac{\partial CP_{\theta}(z_{1},n_{2})}{\partial n_{2}}|_{n_{2}=n_{2}^{*}(z_{1})}\,.$
By this construction, the pre-specified SSR rule $n_{2}^{*}(z_{1})$ satisfies the first-order necessary condition for the optimization problem where the cost function is $\gamma^{*}(z_{1})$. This satisfies the optimality equation and therefore proves the existence of a cost function that is consistent with the given rule.
We then show that $n_{2}^{*}(z_{1})$ is truly the solution to the optimization problem, we will demonstrate that the first-order condition is not just necessary but also sufficient for a maximum, and that this maximum is unique. For the objective function, $CP_{\theta}(z_{1},n_{2})-\gamma(z_{1})n_{2}$. We assume that $CP_{\theta}(z_{1},n_{2})$ is a strictly concave function of $n_{2}$ for $z_{1}\in\mathcal{P}$. Further, $\gamma(z_{1})n_{2}$ is a linear function of $n_{2}$. The sum of a strictly concave function and a linear function is itself strictly concave. Therefore, the objective function is is strictly concave in $n_{2}$ for any given $z_{1}$ and its corresponding $\gamma(z_{1})$.
For strictly concave functions, the first-order condition is both necessary and sufficient for a unique global maximum. This means that the rule $n_{2}^{*}(z_{1})$ is not merely a stationary point; it is guaranteed to be the unique rule that maximizes the objective function. ∎
@@
For the purpose of comparing SSR rules, our primary interest lies in their behavior within the promising zone. Hence, it suffices to establish a one-to-one correspondence between SSR rules and cost functions restricted to this region. Accordingly, it is sufficient for Theorem 1 that the objective function be strictly concave with respective to $n_{2}$ for $z_{1}$ within the promising zone.
@@
The ability to reverse-engineer a SSR rule to find its implied cost function provides a powerful auditing tool for evaluating proposed or existing SSR designs. An analysis might reveal that a seemingly sensible rule implies a $\gamma(z_{1})$ that is erratic, or counter-intuitive. For instance, a rule might implicitly value an additional participant more highly in a moderately promising zone than in a very promising zone, a behavior that may be difficult to justify rationally. Identifying such anomalies can serve as a critical red flag, indicating that the rule could lead to inefficient or logically inconsistent resource allocation under certain interim outcomes. This analytical capability allows for a deeper and more critical appraisal of SSR strategies, moving beyond surface-level performance metrics like average power and sample size to scrutinize the logical coherence of the underlying decision-making process.

### v2

@@ 2 Background of the JT Framework
We consider a clinical trial with a continuous endpoint, where measurements for the treatment and control arms are denoted by $Y_{i}^{trt}$ and $Y_{i}^{ctl}$, $i=1,2,\dots$, respectively. Assuming normality with a common known variance $\sigma^{2}$, we have:
@@
The treatment effect $\theta$ is defined as the difference in means $\theta=\mu_{t}-\mu_{c}$. An interim analysis is planned after collecting $n_{1}$ measurements with a 1:1 allocation ratio. The observed treatment effect is $\hat{\theta}_{1}=\bar{Y}^{trt}-\bar{Y}^{ctl}$ and the corresponding standardized test statistic is denoted by $Z_{1}$.
The final decision uses an inverse-normal combination test with weights $w_{1},w_{2}>0$ satisfying $w_{1}^{2}+w_{2}^{2}=1$ and critical value $C_{\mathrm{crit}}$. Conditional on $Z_{1}=z_{1}$ and total final sample size $n_{2}\geq n_{1}$, the conditional power under effect size $\theta$ can be written in the standard form
$CP_{\theta}(z_{1},n_{2})\;=\;\Phi\!\Big(A(z_{1},n_{2})\Big),\quad A(z_{1},n_{2})\;=\;K\sqrt{n_{2}-n_{1}}\;-\;c(z_{1}),$ (2.1)
with $K=\theta/(2\sigma)$ and $c(z_{1})=(C_{\mathrm{crit}}-w_{1}z_{1})/w_{2}$.
@@
$n_{2}^{JT}(z_{1})=\max_{n_{2}\in[n_{\min},N_{\max}]}\big\{CP_{\theta}(z_{1},n_{2})-\gamma\,(n_{2}-n_{\min})\big\},\,$, (2.2)
where $\gamma>0$ is a fixed cost parameter and $n_{\min},N_{\max}]$ is the allowable range of total sample sizes. The term $CP_{{\theta}}(z_{1},n_{2}(z_{1}))$ represents the conditional power, given $Z_{1}=z_{1}$, evaluated under a prespecified effect size of interest, $\theta$}. The quantity $\gamma$ serves as a tuning parameter that represents the acceptable “rate of exchange” between conditional power and sample size. It can be tied to the cost of additional subjects and is used to control the extent to which the sample size may be increased based on the interim data.
@@ 3 Optimal Sample Size Re-estimation with Dynamic Cost Functions
$\arg\max_{n_{2}\in[n_{\min},N_{\max}]}\left\{CP_{\theta}(z_{1},n_{2})-\gamma(z_{1})n_{2}\right\}\,.$ (3.1)
@@ 3.1 Likelihood-ratio based $\gamma(z_{1})$
$\max_{n_{2}\in\mathcal{N}}P_{\theta}(\mathrm{Reject}\,H_{0}|\Theta_{1})\hskip 14.22636pt\text{subject to}\hskip 14.22636ptE_{0}[n_{2}(Z_{1})]\leq B_{0},\hskip 14.22636pt\text{and}\hskip 14.22636ptE_{\theta}[n_{2}(Z_{1})]\leq B_{\theta}\,$, (3.2)
@@
We let $f_{0}(z_{1})$ and $f_{\theta}(z_{1})$ denote densities of $Z_{1}$ under $\Theta_{0}$ and $\Theta_{1}$, respectively. We assume that $f_{\theta}(z_{1})>0$ for $z_{1}\in\mathbb{R}$. For nonnegative Lagrange multipliers $\lambda_{1},\lambda_{2}\geq 0$, the Lagrangian is
$\mathcal{L}\!\left(n_{2}(\cdot),\lambda_{1},\lambda_{2}\right)=\int\Big\{CP_{\theta}\!\big(z_{1},n_{2}(z_{1})\big)\,f_{\theta}(z_{1})-\big(\lambda_{1}f_{0}(z_{1})+\lambda_{2}f_{\theta}(z_{1})\big)\,n_{2}(z_{1})\Big\}\,\mathrm{d}z_{1}.$
Because the integral is separable in $z_{1}$, that is, $n_{2}$ enters the objective only through its value at the same $z_{1}$ and there are no cross $z_{1}$ couplings (no terms involving $n_{2}(z_{1}^{\prime})$ with $z_{1}^{\prime}\neq z_{1}$), maximizing $\mathcal{L}$ over $n_{2}(\cdot)$ reduces to pointwise maximization of the integrand. On the set where $f_{\theta}(z_{1})>0$, dividing by $f_{\theta}(z_{1})$ does not change the maximizer and yields the equivalent problem
$\max_{\,n_{2}(z_{1})}\;\Big\{CP_{\theta}\big(z_{1},n_{2}(z_{1})\big)-\Big(\lambda_{2}+\lambda_{1}\frac{f_{0}(z_{1})}{f_{\theta}(z_{1})}\Big)\,n_{2}(z_{1})\Big\},$
i.e., a pointwise trade-off between conditional power and a dynamic cost term proportional to the likelihood ratio $f_{0}(z_{1})/f_{\theta}(z_{1})$. We denote the solution of this optimization problem by $n_{2}^{LR}(z_{1})$. This formulation yields a dynamic cost function in the JT fashion, where the cost is given by $\gamma(z_{1})=\lambda_{2}+\lambda_{1}\frac{f_{0}(z_{1})}{f_{\theta}(z_{1})}$. The term $\frac{f_{0}(z_{1})}{f_{\theta}(z_{1})}$ is the likelihood ratio. A large value of this ratio indicates that the observed interim statistic $z_{1}$ is more likely to arise under the null hypothesis, thereby increasing the cost of adding a sample and making it harder to justify a sample size increase.
3.1.1 Comparing $n_{2}^{JT}$ with $n_{2}^{LR}$
Let $n_{2}^{\mathrm{JT}}:\mathbb{R}\to\mathcal{N}$ denote the JT rule and $n_{2}^{\mathrm{LR}}:\mathbb{R}\to\mathcal{N}$ the proposed likelihood ratio based rule; we write $n_{2}^{\mathrm{JT}}(z_{1})$ and $n_{2}^{\mathrm{LR}}(z_{1})$ for their values at a given $z_{1}$. For a fixed cost parameter $\gamma$, we write $n_{2}^{\mathrm{JT}}(\cdot;\gamma)$. We now formally compare the performance of the JT rule, $n_{2}^{JT}$, with our proposed likelihood-ratio based rule, $n_{2}^{LR}$. For a fixed cost parameter $\gamma$, let $n_{2}^{JT}(\cdot,\gamma)$ be the optimal rule from the JT method. We claim that $n_{2}^{JT}(\cdot,\gamma)$ is not the optimal solution for the following constrained optimization problem:
$\min_{n_{2}}E_{0}[n_{2}(Z_{1})]\quad\text{subject to}\quad P_{\theta}(\mathrm{Reject}\,H_{0};n_{2})=P_{\theta}(\mathrm{Reject}\,H_{0};n_{2}^{JT}(\cdot,\gamma))\quad\text{and}\quad E_{\theta}[n_{2}(Z_{1})]=E_{\theta}[n_{2}^{JT}(Z_{1},\gamma)]\,.$
@@ 3.1.1 Comparing $n_{2}^{JT}$ with $n_{2}^{LR}$
$\mathcal{L}(n_{2},\mu_{1},\mu_{2})$ $=E_{0}[n_{2}(Z_{1})]-\mu_{1}\left(P_{\theta}(\mathrm{Reject}\,H_{0};n_{2})-P_{\theta}(\mathrm{Reject}\,H_{0};n_{2}^{JT}(\cdot,\gamma))\right)$
@@
For a function $n_{2}$ to be a minimizer of $\mathcal{L}(n_{2},\mu_{1},\mu_{2})$, the integrand must be minimized for each $z_{1}$. We treat $n_{2}$ as a continuous variable for the purpose of optimization, a standard approach for this class of problems. While the sample size must be an integer, treating it as a continuous variable allows for the use of calculus-based optimization. The resulting continuous solution can then be rounded to the nearest integer. The performance loss from this approximation is typically negligible. Thus, a necessary condition for $n_{2}$ to be a minimizer is:
@@
For the JT method, the corresponding condition is $\frac{\partial CP_{\theta}(z_{1},n_{2}(z_{1}))}{\partial n_{2}(z_{1})}=\gamma$, which is a constant. Since the right-hand side of our derived condition is a function of $z_{1}$, this proves that $n_{2}^{JT}(z_{1})$ is not the rule that minimizes $E_{0}[N]$ for its given power and average sample size under the alternative hypothesis.
@@
3.1.2 Pareto optimality of $n_{2}^{LR}$
The likelihood-ratio based re-estimation rule, $n_{2}^{LR}$ is a solution to multi-objective optimization problem. As such, it possesses a fundamental property known as Pareto optimality.
@@ 3.1.2 Pareto optimality of $n_{2}^{LR}$
A rule $n_{2}^{*}$ is Pareto optimal if no other rule $n_{2}$ exists that is strictly better on at least one objective while being no worse on the others.
@@
For $\lambda_{1}\geq 0$ and $\lambda_{2}\geq 0$, $n_{2}^{LR}$ maximizes $\mathcal{L}(n_{2})=P_{\theta}(\mathrm{Reject}\,H_{0};n_{2})-\lambda_{2}1E_{0}[n_{2}(Z_{1})]-\lambda_{2}E_{\theta}[n_{2}(Z_{1})]$. Assume that $n_{2}^{LR}$ is not Pareto optimal. By definition, this means there exists a rule $n^{*}_{2}$ such that
  1. $P_{\theta}(\mathrm{Reject}\,H_{0};n_{2}^{*})>P_{\theta}(\mathrm{Reject}\,H_{0};n^{LR}_{2})$  ,
@@ 2. $E_{0}[n^{*}_{2}(Z_{1})]\leq E_{0}[n^{LR}_{2}(Z_{1})]$  ,
Then, we got $\mathcal{L}(n^{*}_{2})>\mathcal{L}(n^{LR}_{2})$, which contradicts to the fact that $n_{2}^{LR}$ maximizes $\mathcal{L}(n_{2})$. ∎
The Pareto optimality of $n_{2}^{LR}$ establishes it as an efficient and well-balanced solution for the multi-objective problem of maximizing power while controlling the average sample size under both $\Theta_{0}$ and $\Theta_{1}$.
In the context of sample size re-estimation, Pareto optimality means that the likelihood-ratio based rule avoids wasteful designs. If one attempts to reduce sample size under the null, the only way to do so is by lowering power; conversely, if one increases power, additional participants must be recruited. Thus, $n_{2}^{LR}$ represents an efficiency frontier: it offers the best achievable balance between power and resource use. This provides reassurance to practitioners that adopting $n_{2}^{LR}$ does not leave potential efficiency gains unexploited.
@@ 3.2 Return on investiment (ROI) based $\gamma(z_{1})$
Let $p(z_{1})=P(\text{effective}|z_{1})$ be the posterior probability of the treatment being effective, and $CP_{\theta}(z_{1},n_{2}(z_{1}))$ be the conditional power. The expected net return can then be expressed as:
@@ 4 Representation Theorem for Sample Size Re-estimation Rules
Let $\mathcal{P}\subseteq\mathbb{R}$ denote the promising zone and let $\mathcal{N}:=(n_{\min},N_{\max})$. Suppose that for every $z_{1}\in\mathcal{P}$ the map $n_{2}\mapsto CP_{\theta}(z_{1},n_{2})$ is differentiable and strictly concave on $\mathcal{N}$. Let $n_{2}^{*}:\mathcal{P}\to\mathcal{N}$ be an SSR rule taking values in the interior. Then there exists a unique function
$\gamma^{*}:\mathcal{P}\to\mathbb{R},\qquad\gamma^{*}(z_{1}):=\left.\frac{\partial}{\partial n_{2}}CP_{\theta}(z_{1},n_{2})\right|_{n_{2}=n_{2}^{*}(z_{1})},$
such that, for every $z_{1}\in\mathcal{P}$, the point $n_{2}^{*}(z_{1})$ is the unique maximizer of
$\max_{n_{2}\in\mathcal{N}}\;\Big\{CP_{\theta}(z_{1},n_{2})-\gamma^{*}(z_{1})\,n_{2}\Big\}.$
Moreover, if $\tilde{\gamma}:\mathcal{P}\to\mathbb{R}$ satisfies $n_{2}^{*}(z_{1})\in\arg\max_{n_{2}\in\mathcal{N}}\{CP_{\theta}(z_{1},n_{2})-\tilde{\gamma}(z_{1})n_{2}\}$ for all $z_{1}\in\mathcal{P}$, then $\tilde{\gamma}=\gamma^{*}$ on $\mathcal{P}$.
@@
Fix any $z_{1}\in\mathcal{P}$ and define
$\gamma^{*}(z_{1}):=\left.\frac{\partial}{\partial n_{2}}CP_{\theta}(z_{1},n_{2})\right|_{n_{2}=n_{2}^{*}(z_{1})}.$
Consider the one dimensional objective
$\phi_{z_{1}}(n_{2})\ :=\ CP_{\theta}(z_{1},n_{2})-\gamma^{*}(z_{1})\,n_{2},\qquad n_{2}\in\mathcal{N}.$
By assumption, $n_{2}\mapsto CP_{\theta}(z_{1},n_{2})$ is strictly concave and differentiable on $\mathcal{N}$; subtracting a linear term preserves strict concavity. Hence $\phi_{z_{1}}$ is strictly concave on $\mathcal{N}$. Moreover,
$\frac{\partial}{\partial_{n_{2}}}\phi_{z_{1}}(n_{2})\ =\ \frac{\partial}{\partial_{n_{2}}}CP_{\theta}(z_{1},n_{2})-\gamma^{*}(z_{1}),$
so $\frac{\partial}{\partial_{n_{2}}}\phi_{z_{1}}\big(n_{2}^{*}(z_{1})\big)=0$ by construction. For a strictly concave, differentiable objective, the first order condition is both necessary and sufficient and the maximizer is unique; thus $n_{2}^{*}(z_{1})$ is the unique maximizer of $\phi_{z_{1}}$ over $\mathcal{N}$.
For uniqueness of the cost, suppose $\tilde{\gamma}:\mathcal{P}\to\mathbb{R}$ also satisfies $n_{2}^{*}(z_{1})\in\arg\max_{n_{2}\in\mathcal{N}}\{CP_{\theta}(z_{1},n_{2})-\tilde{\gamma}(z_{1})n_{2}\}$ for all $z_{1}\in\mathcal{P}$. Strict concavity implies uniqueness of the maximizer and, hence, the first order condition at $n_{2}^{*}(z_{1})$: $\frac{\partial}{\partial_{n_{2}}}CP_{\theta}(z_{1},n_{2}^{*}(z_{1}))=\tilde{\gamma}(z_{1})$. By the definition of $\gamma^{*}$, we conclude $\tilde{\gamma}(z_{1})=\gamma^{*}(z_{1})$ for all $z_{1}\in\mathcal{P}$. ∎
@@
For the purpose of comparing SSR rules, our primary interest lies in their behavior within the promising zone. Hence, it suffices to establish a one-to-one correspondence between SSR rules and cost functions restricted to this region. Accordingly, it is sufficient for Theorem 1 that the objective function be strictly concave with respect to $n_{2}$ for $z_{1}$ within the promising zone.
@@
The ability to reverse-engineer a SSR rule to find its implied cost function provides a powerful auditing tool for evaluating proposed or existing SSR designs. An analysis might reveal that a seemingly sensible rule implies a $\gamma$ function that is erratic, or counter-intuitive. For instance, a rule might implicitly value an additional participant more highly in a moderately promising zone than in a very promising zone, a behavior that may be difficult to justify rationally. Identifying such anomalies can serve as a critical red flag, indicating that the rule could lead to inefficient or logically inconsistent resource allocation under certain interim outcomes. This analytical capability allows for a deeper and more critical appraisal of SSR strategies, moving beyond surface-level performance metrics like average power and sample size to scrutinize the logical coherence of the underlying decision-making process.
Beyond retrospective evaluation, the representation theorem offers a constructive paradigm: sponsors may specify a rational cost function that reflects their risk-benefit preferences and derive the corresponding SSR rule. This approach enhances transparency and provides regulators with a principled justification that goes beyond operating characteristics to articulate the rationale underlying the design.